OpenStudy (anonymous):

A 12-foot ladder is placed 4 feet from the wall. Approximately how far up the wall will the ladder reach? 11 16 8

Parth (parthkohli):



Parth (parthkohli):

\( \color{Black}{\Rightarrow 4^2 + ?^2 = 12^2}\) \( \color{Black}{\Rightarrow ?^2 = 144 - 16}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow ? = \sqrt{128} = \dots }\)

OpenStudy (anonymous):

11

Parth (parthkohli):

Yeah